OpenStudy (anonymous):

Can somebody help me fill out a punnett square or tech me how to do it PLEASE

OpenStudy (anonymous):

In pea plants, purple flowers (P) are dominant over white flowers (p) and tall stems (T) are dominant over short stems (t). Complete a dihybrid cross for parents with the genotypes: PpTT × ppTt and answer the following in complete sentences. Describe how you would set up a Punnett square for this cross. List the likelihood of each possible offspring genotype. List the likelihood of each possible offspring phenotype.

OpenStudy (anonymous):

Im trying to slove this problem and so far i have

OpenStudy (anonymous):

1. you set up 16 boxes for a punnett square and the you fill you the alleles on top or the left side Both parents produce 75% each of PT PT pT pT. 2. PpTT ppTt PT PT pT pT x pT pt pT pt

OpenStudy (anonymous):

?

OpenStudy (anonymous):

1. you set up 16 boxes for a punnett square and the you fill you the alleles on top or the left side Both parents produce 75% each of PpTT × ppTt 2. PpTT ppTt PT PT pT pT x pT pt pT pt Phenotypes PT PT pT pT PpTT PpTT ppTT ppTT PpTt PptT ppTt ppTt PpTT PpTT ppTT ppTT PpTt PptT ppTt pptT pT pt pT pt PpTT = 4/16 ppTT = 4/16 PpTt = 2/16 PptT = 2/16 pptT = 3/16 pptT = 1/16